Diffusion Morning Report 04/28/06

Ted Berg tberg at snf.stanford.edu
Fri Apr 28 06:40:37 PDT 2006


Good Morning All,
    Looking pretty good this morning. Burn box cleaned this morning. Epi 
is back to Yellow. Everything else looks OK for now. Laser
elipsometer went back to NJ for repair.

Diffusion 	  	  	  	  	 
  	Date: 	4/28/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	Up 	  	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	19.5/30 	  	 
Tylan 8-SiGe 	Up 	  	28.46/30 	  	 
Tylan 9-Poly 	Up 	  	2.39/30 	  	 
Tylan 10-Nitride 	Up 	  	23.46/30 	  	 
Tylan 11-BPSG-LTO 	up 	low dep rate-due to boat 	3.5/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Down 	Testing 	  	  	 
  	  	  	  	  	 
Epi 	yellow 	possible higher leak rate 	  	  	 
  	  	  	  	  	 
Burn Box 	Up 	cleaned this morning 	  	  	 
ZYGO 	up 	  	   	  	 
  	  	  	  	  	 
GAAS 22 	UP 	  	  	  	 
GAAS 23 	up 	  	  	  	 
  	  	  	  	  	 
SiH4 for furnaces 	  	A on line 590 psi 	 changeover working well  	  	 




-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060428/b549b372/attachment.html>


More information about the diffusion mailing list